An analysis of the Business School graduates found that 211 out of 310 randomly selected graduates used a statistical inference technique during their first year of emplyment. (a) Calculate a 90% confidence interval for the proportion of graduates who used a statistical inference technique within the first year of their employment. Present your work. (b) Interpret the confidence interval. (c) If a new follow-up study is to be undertaken, what sample size should be taken in order to estimate the true proportion of graduates who used a statistical inference technique in the first year of employment within 2%, with a confidence interval of 90%? (d) If a new follow-up study is to be undertaken, what sample size should be taken in order to estimate the true proportion of graduates who used a statistical inference technique in the first year of employment within 2%, with a confidence interval of 99%?
Answer #1

Solution:-

a) 90% confidence interval for the proportion of graduates who used statistical inference technique within first year of employment is C.I = ( 0.6371, 0.7242).

p(1 - p)

C.I = 0.68065 + 1.645*0.02648

C.I = 0.68065 + 0.04356

C.I = ( 0.6371, 0.7242)

b) If repeated samples were taken and the 90% confidence interval was computed for each sample, 90% of the intervals would contain the population proportion.

c) The required sample size is 1471.

M.E = 0.02, C.L = 90%

p(1 - p)

n =left ( rac{z_{alpha /2}}{M.E} ight )^{2} imes p(1-p)

1.645 2 n-( 0.02 × 0.68065(1-0.68065) rl

n = 1470.51

n = 1471

d) The required sample size is 3607.

M.E = 0.02, C.L = 99%

p(1 - p)

n =left ( rac{z_{alpha /2}}{M.E} ight )^{2} imes p(1-p)

= (2.5 ),068065(1-0.68065) n-( 0,05) ×0.68065(1-0.68065) rl

n = 3606.03

n = 3607